### v1

# AB-Training: A Communication-Efficient Approach for Distributed Low-Rank Learning

Abstract.
Communication bottlenecks hinder the scalability of distributed neural network training, particularly on distributed-memory computing clusters. To significantly reduce this communication overhead, we introduce AB-training, a novel data-parallel training method that decomposes weight matrices into low-rank representations and utilizes independent group-based training. This approach consistently reduces network traffic by 50% across multiple scaling scenarios, increasing the training potential on communication-constrained systems. Our method exhibits regularization effects at smaller scales, leading to improved generalization for models like VGG16, while achieving a remarkable 44.14 : 1 compression ratio during training on CIFAR-10 and maintaining competitive accuracy. Albeit promising, our experiments reveal that large batch effects remain a challenge even in low-rank training regimes.
Distributed training, Low-rank representation, Neural networks, Communication efficiency
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 53rd International Conference on Parallel Processing; August 12–15, 2024; Gotland, Swedenccs: Computing methodologies Neural networksccs: Computing methodologies Machine learning algorithmsccs: Networks Network performance analysisccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Batch learning
Figure 1. Average total interconnect traffic measured during the training of a Vision Transformer (ViT) B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ with traditional, synchronous, full-rank data parallel training as implemented by PyTorch’s DistributedDataParallel (DDP) class and our AB training method. The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
[…]
Average total interconnect traffic measured during the training of a Vision Transformer B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$. The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
1. Introduction
The relentless pursuit of predictive performance in machine learning has led to neural networks of unprecedented scale and complexity. These massive state-of-the-art models excel in tasks ranging from image classification and natural language processing to scientific simulations. However, this performance comes at a steep computational cost, as training large models often demands large datasets. These datasets often exceed the capabilities of individual devices, necessitating distributed training approaches.
Data parallelism (DP) is the most common distributed training technique for neural networks (Ben-Nun and Hoefler, 2019). It replicates the model across multiple devices, each of which processes a subset of the data, known as the local batch, before aggregating the gradients and updating model parameters. A persistent challenge in data parallel (DP) training is the communication bottleneck created by synchronizing large model representations across compute nodes (Niu et al., 2011). This bottleneck severely limits scalability, particularly on distributed-memory computing systems with potential network constraints. Large batch effects (Goyal et al., 2018), pose an additional challenge for maintaining predictive performance, as accuracy degrades once the global batch size crosses a critical threshold.
To address these challenges, we propose a novel low-rank training method based on the singular value decomposition (SVD) and a hierarchical training scheme designed to significantly reduce communication overheads in distributed settings. Our method leverages the observation that low-rank representations can promote regularization (Coquelin et al., 2024) and minimize the amount of data transferred. Key contributions of our work include:
  • We find a 50% reduction in network traffic during distributed training on average as compared to traditional synchronous DP training, enabling experimentation with larger batch sizes and potentially accelerating training within distributed computing environments without high-speed interconnects.
  • Our method demonstrates improved regularization effects at smaller scales, particularly on the Vision Transformer (ViT), leading to better predictive performance on unseen data. However, it also sheds light onto the challenge of reducing large batch effects without hyperparameter tuning.
  • In an ideal scenario, AB training achieves a compression ratio of 44.14 : 1 for the VGG16 (Liu and Deng, 2015) model when trained on CIFAR-10 (Krizhevsky, 2009) while maintaining a competitive accuracy. In more standard scenarios, it achieved compression values ranging from 1.19 : 1 to 2.54 : 1 when exceeding the accuracy of traditional DP training.
Our work highlights promising directions for how to fully harness the potential of low-rank training in massively scaled scenarios, with implications for both distributed-memory computing and the broader ML community.
2. Related Work
2.1. Distributed Training of Neural Networks
The computational demands of training large neural networks have driven the development of distributed training paradigms, with data parallelism as a cornerstone approach. In a traditional data parallel (DP) training scheme, the model is replicated across multiple devices, each processing a subset of the training data independently. The data subset that one model instance processes is the local batch, while the global batch is the set of all local batches. After each forward-backward pass, gradients are aggregated across all model instances before the model is updated by the optimizer. As the number of compute resources and the model size increase, communication overheads incurred during gradient synchronization become a significant constraint, particularly in systems without high-bandwidth interconnects. Other parallelization strategies, such as model parallelism, introduce alternative ways to distribute training across devices, but come with their own communication and load-balancing complexities.
Several techniques have been proposed to mitigate the communication bottleneck in data parallel training, including gradient accumulation, topology-aware communication patterns, asynchronous methods, and gradient compression (Vogels et al., 2019; Abrahamyan et al., 2022). Methods like H-SGD (Lin et al., 2020b) and DASO (Coquelin et al., 2022) utilize localized synchronization within smaller groups between global updates, potentially reducing communication costs by leveraging the network’s topology. Asynchronous approaches typically make use of a parameter server for aggregating gradients (Niu et al., 2011) and often require extensive hyperparameter tuning. Despite these advancements, the scalability of data parallelism remains constrained by communication bottlenecks and large batch effects, where increasing the global batch size negatively impacts generalization performance.
2.2. Low-Rank Neural Network Training
A promising avenue for reducing both the memory footprint and computational cost of neural networks lies in low-rank training. Singular Value Decomposition (SVD) plays a crucial role in this domain. By factoring a weight matrix $W$ into three matrices, $U$, $\Sigma$, and $V^{T}$, and retaining only the $k$ largest singular values, a low-rank approximation with significantly fewer parameters can be obtained given a sufficiently large $k$ value. Note that aggressive compression can lead to a loss of approximation quality.
Several methods leverage SVD for compressing neural networks during training. Some train directly on the $U$, $\Sigma$, and $V$ matrices (Schotthöfer et al., 2022; Coquelin et al., 2024), while others create different representations using the SVD and training on them (Liebenwein et al., 2021; Yang et al., 2020a). Both approaches can offer regularization-like benefits, potentially improving generalization (Winata et al., 2020; Cahyawijaya et al., 2021; Phan et al., 2020).
Gradient compression using SVD has found application in federated learning scenarios (Wang et al., 2023b), where bandwidth constraints necessitate reduced communication. While these techniques address computational challenges in neural network training, their integration within distributed environments to specifically optimize communication remains an active area of research.
Low-rank factorizations can heavily reduce the communication required during distributed training. Existing approaches simply train low-rank models in parallel, but do not make any efforts to use the distributed system to their advantage (Coquelin et al., 2024; Wang et al., 2023a, 2021). To the best of our knowledge, there exist no methods that leverage both distributed architectures and low-rank representations to train neural networks.
3. Methodology
Data parallel training of neural networks is constrained by the the communication bottleneck caused by synchronizing massive model representations across compute nodes and large batch effects, a phenomenon where accuracy degrades at large global batch sizes. In an attempt to tackle these challenges, our proposed training method integrates low-rank weight representations within a data parallel framework. It leverages the reduced network capacity stemming from topologically informed communication patterns and low-rank training while escaping local minima using group training.
We utilize SVD to decompose weight matrices into low-rank representations, significantly reducing the amount of data communicated during training. We decompose the weights of a given layer as
(1) $W_{m\times n}=A_{m\times k}B_{k\times n}=\left(U_{m\times k}\Sigma_{k\times k}^{1/2}\right)\left(\Sigma_{k\times k}^{1/2}V_{k\times n}^{\top}\right)$
where $U$, $\Sigma$, and $V$ are determined by the SVD of $W$ and $k<\texttt{min}\left(m,n\right)$. Using a low-rank representation in training has the added benefit of reducing overfitting, a common problem in training neural networks (Yang et al., 2020b; Hu et al., 2021).
Large batch effects can be trickier to deal with. Consider the standard batch stochastic gradient descent (SGD) optimization step
(2) $W_{t}=W_{t-1}-\frac{\eta}{N}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{t-1}\right),$
where $W_{t}$ is a parameter at time $t$ in training, $\eta$ is the learning rate, $N$ is the batch size, and $Q_{i}\left(W\right)$ is the value of the loss function for the $i$-th data element. It is easily shown that for a given iteration $k>0$, the weights are simply the sum of all the previous update steps or
(3) $W_{k}=W_{0}-\frac{\eta}{B}\sum_{s=1}^{k-1}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{s}\right)$
With a larger batch size, $B$, the magnitude of the gradients shrinks. This removes the randomness inherent to SGD’s success and makes the network more prone to getting stuck in local minima (Chaudhari et al., 2019).
To mitigate this effect, we draw motivation from ensemble approaches. By dividing the DP model replicas into subgroups to be trained on independent data subsets, we essentially form an ensemble of models. Individually, these groups converge towards different local minima compared to full-system training, promoting exploration of the loss landscape.
After training in subgroups, we average the independently trained models. This merging strategy has seen some success in federated learning methods (Giuseppi et al., 2022). Compared to federated use-cases, which typically deal with models trained independently on biased datasets, AB training benefits from randomly sampling from the same dataset and thus has unbiased batches during our independent training phases.
During the independent training phases, each group trains on either $A$ or $B$, while the other component remains fixed. As the orthogonal component of $W$ is stable later in training (Coquelin et al., 2024), we only train on one of $A$ or $B$ in an attempt to reduce the divergence between the independent groups. At the end of the independent training phase, we average $A$ and $B$ across all workers to merge the independently trained models together.
Our AB training procedure consists of the following phases:
Distributed warmup:
Initial training with full-rank data parallelism allows the weights to adapt rapidly away from their random initialization. To avoid early divergence, we employ a traditional, full-rank data parallel warmup phase for a given number of iterations (warmupSteps).
Independent AB decomposition:
Each model instance independently computes an AB decomposition of its weight matrices as per Equation 1. This AB decomposition allows us to represent the weight matrices more compactly, reducing the amount of data to be communicated during training. The rank of this approximation is determined by a user-provided hyperparameter (sigmaCutoff) times the largest singular value. Values smaller than this are removed.
Group-only training:
Half of the independent groups will train the $A$ matrix, denoted as the $A$ groups, while the other groups will train $B$, the $B$ groups. For a given number of iterations (numABSteps), processes in the $A$ group only train the $A$ matrices, while processes in the $B$ group only train the $B$ matrices, the matrix not being trained remains frozen in this phase.
Synchronization and update:
We average both the trained and frozen $A$ and $B$ matrices independently to synchronize the model instances across all processes. This is crucial for ensuring that the overall model remains consistent across all compute nodes. Afterwards, we reconstruct the full-rank weight matrices from the updated $A$ and $B$ components.
Full-rank rebound:
We then train the full-rank network with traditional data parallelism for a short period, which we refer to as the fullRankReboundSteps. This phase promotes convergence and helps mitigate potential accuracy degradation observed in low-rank training scenarios.
A formalized algorithm is shown in Algorithm 1 and the method diagram is shown in Figure 2.
Figure 2. The AB training procedure.
[…]
The AB training procedure. See pseudocode or description in text.
Input : Model, $M$ , training data, $numTrainingSteps$, hyperparameters ($warmupSteps$, $numABSteps$, $fullRankReboundSteps$)
1
2if procId $\leq$ worldSize / 2 then
3      $groupTrainTarget\leftarrow A$
4else
5      $groupTrainTarget\leftarrow B$
6
7for $i\leftarrow 1$ to warmupSteps do
8       traditionalFullRankDPTraining()
9repeat
10       foreach $W$ in $M$ do
11             $A,B\leftarrow\texttt{abDecomposition}(W)$triangleright Equation 1
12             removeSmallSingularValues()
13      startLearningRateRebound()
14       $groupTrainTarget$.setTrainable(True)
15
16      for $i\leftarrow 1$ to numABSteps do
17             independentSubgroupTraining()
18      $A\leftarrow\texttt{allReduce}(A/worldSize$)
19       $B\leftarrow\texttt{allReduce}(B/worldSize$)
20       foreach $W$ in $M$ do
21             $W\leftarrow AB$
22      startLearningRateRebound()
23       for $i\leftarrow 1$ to fullRankReboundSteps do
24             traditionalFullRankDPTraining()
25
26until numTrainingSteps = completed steps
Algorithm 1 The AB training method. $W$ is a parameter of the input model $M$ and $worldSize$ is the number of workers used for traditional DP training.
Popular neural network optimizers like AdamW (L. and H., 2018) rely on second-order derivative approximations. When we change the network’s structure (e.g., between low-rank and full-rank representations), these optimizer states become incompatible due to their dimensionally. To address this, we use a “learning rate rebound” strategy. This involves reducing the learning rate to near-zero and gradually increasing it back to its scheduled value over a user-provided number of steps, allowing the optimizer to adjust to the new parameter representations. We do a learning rate rebound after each time the network’s parameters change shape.
4. Experimental Evaluation
To evaluate the effectiveness of our AB training approach, we conducted experiments on established neural network architectures. Our primary objectives were to demonstrate significant reductions in communication overhead, potential regularization benefits, scalability, and achievable compression during training. We focused on the ResNet-50 (He et al., 2016) and Vision Transformer (Dosovitskiy et al., 2021) B/16 models trained on the ImageNet-2012 dataset (Krizhevsky et al., 2012) for image classification as well as VGG16 (Liu and Deng, 2015) trained on CIFAR-10 (Krizhevsky, 2009). We compare our results against traditional DP training as implemented by PyTorch’s DDP package, referred to as ‘Traditional DDP’ or ‘Trad. DDP’, using the same hyperparameters (HPs). We also show results for the AB training setup without any individual trained groups which are denoted as ‘No Group’. In these experiments, the $B$ matrix is trained using traditional DP training in place of the full-rank weights. The ViT HPs are those used in the original source, while the ResNet-50 parameters utilize the same learning rate scheduler and the default values for the AdamW (L. and H., 2018) optimizer. The VGG16 HPs are as shown in (Coquelin et al., 2024).
We analyze our training method using two scaling strategies: constant local batch size scaling and constant global batch size scaling. In constant local batch size scaling, the global batch size increases with the number of GPUs. In constant global batch size scaling, we maintain a constant global batch size while proportionally decreasing the local batch size as we increase the number of GPUs. This emphasizes the communication overhead reductions achieved by our method. Measurements represent the average of three runs, each of which utilizes a different random seed. Models are initialized using orthogonal initialization (Hu et al., 2020).
4.1. Computational Environment
[…]
4.2. Datasets and Models
We used the ImageNet-2012 dataset, which contains 1.2 million images, for our scaling experiments. Basic image augmentation was applied, including normalization, random resizing, random cropping, and a random flip. We trained two models on this dataset: ResNet-50 (He et al., 2016) and Vision Transformer B/16 (ViT) (Dosovitskiy et al., 2021), chosen for their widespread use and distinct architectures.
We also trained VGG16 on the CIFAR-10 dataset, which contains 50,000 images. Basic image augmentation was applied in the same way as for the ImageNet experiments.
All models were trained with the AdamW (Loshchilov and Hutter, 2019) optimizer. Other training hyperparameters can be found in Appendix 1.
4.3. Hyperparameter Considerations
Careful hyperparameter tuning is essential for balancing the communication reduction achieved through our method with potential impacts on training stability and accuracy. Key hyperparameters (HPs) include the warmup and full-rank rebound phase durations, the number of AB training iterations, the AB decomposition rank-reduction parameter, and the frequency of SVD and synchronization steps. Additionally, the latter will influence the trade-offs between regularization and communication efficiency.
Our hyperparameter search on CIFAR-100 (Krizhevsky, 2009) with a minified Vision Transformer model (patch size of eight instead of 16, six heads instead of twelve, depth of six instead of twelve) using Propulate (Taubert et al., 2023) suggested the following guidelines:
  • warmupSteps: 20% of the total training steps
  • numABSteps: 3.33% of the total training steps
  • fullRankReboundSteps: $0.25\cdot\texttt{numABSteps}$
  • Learning rate rebound steps: $0.5\cdot\texttt{numABSteps}$
4.4. Constant Local Batch Size Scaling
In this experiment, we maintain a fixed local batch size while increasing the available compute resources and decreasing the number of training steps per epoch, resulting in an increasing global batch size but less training iterations. For example, if the global batch size for a two-node (eight-GPU) run is 2,048 and it is trained for 600 iterations in each epoch, the corresponding four-node (16-GPU) run would have a global batch size of 4,096 and be trained for 300 iterations each epoch. This strategy investigates the communication efficiency gains achievable as we increase the available compute resources, even as the potential negative impacts of large batch sizes become more pronounced. It is the most likely method by which a non-expert user will scale their training.
This experiment tests how well AB training reduces communication requirements and if the independent training groups can effectively train networks independently. We expect that if the independent training groups make similar updates, the averaged model will maintain accuracy and exhibit increased compression. However, significant divergence in individual groups could lead to lower accuracy and reduced compression as the individual updates conflict. If this is the case, then AB training without independent groups will begin to outperform AB training with groups.
Figure 3 shows the highest top 1 accuracy achieved by each model during training. Figure 4 shows the lowest binary cross-entropy loss values achieved across all epochs. The compression values for the trained models are shown in Figure 5.
(a) Vision Transformer B/16
(b) ResNet-50
Figure 3. Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. Error are plotted, though not always visible.
[…]
Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2.
(a) Vision Transformer B/16
(b) ResNet-50
Figure 4. Lowest binary cross-entropy (the loss function used during training) measurements during each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. Error are plotted, though not always visible.
[…]
Lowest binary cross-entropy (the loss function used during training) measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. AB training is shown to have a regularization effect for both ResNet-50 and the Vision Transformer. Large batch effects are also visible as the loss values get worse (larger) as the global batch size increases.
Figure 5. Average compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
[…]
Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
4.5. Constant Global Batch Size Scaling
In this experiment, we maintain a fixed global batch size while scaling compute resources. This contrasts with the standard approach of keeping local batch sizes constant, emphasizing the communication overhead reductions achieved by AB training. With a fixed global batch size of 4,096, a four-node run would have a local batch size of 256, whereas an eight-node run would have a local batch size of 128.
This experiment directly quantifies the communication benefits of our method in scenarios where local batch sizes become very small due to scaling. It also provides insight into the ability of independent training groups to learn meaningful representations when the batch size used for independent training is much smaller than the global batch size. Results for these runs are shown in Figures 6 and 7.
(a) Vision Transformer B/16
(b) ResNet-50
Figure 6. Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant global batch size of 4,096. Error are plotted, though not always visible.
[…]
Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
Figure 7. Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
[…]
Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
5. Discussion
Table 1. Average interconnect traffic, in GB/s, on each processes. $N$ represents either the number of nodes for the constant global batch size experiments or the global batch size in thousands for the constant local batch size experiments.
Constant Local
Batch Size
Constant Global
Batch Size
$N$ Baseline AB Baseline AB
ViT 2 $4.67\pm 0.00$ $2.10\pm 0.07$
4 $4.77\pm 0.06$ $2.12\pm 0.01$ $4.81\pm 0.05$ $1.86\pm 0.39$
8 $4.58\pm 0.00$ $1.76\pm 0.29$ $7.05\pm 1.03$ $3.45\pm 0.18$
16 $4.44\pm 0.10$ $1.74\pm 0.07$ $13.66\pm 0.16$ $5.33\pm 0.70$
32 $4.07\pm 0.20$ $1.58\pm 0.24$ $14.09\pm 3.47$ $7.88\pm 0.02$
ResNet-50 2 $2.06\pm 0.01$ $1.02\pm 0.00$
4 $2.07\pm 0.01$ $0.98\pm 0.05$ $2.06\pm 0.01$ $1.00\pm 0.01$
8 $2.00\pm 0.05$ $1.05\pm 0.05$ $2.67\pm 0.05$ $1.16\pm 0.14$
16 $1.96\pm 0.12$ $1.15\pm 0.00$ $3.28\pm 0.25$ $1.39\pm 0.21$
32 $1.62\pm 0.03$ $1.05\pm 0.01$ $3.50\pm 1.00$ $1.76\pm 0.47$
Across both scaling scenarios, we find a consistent reduction of approximately 50% on average network traffic, demonstrating the effectiveness of our AB training method in reducing communication overheads. Comprehensive results are listed in Table 1, which further highlights the significant bandwidth requirements for training even moderately sized neural networks, a challenge that would be exacerbated for larger, state-of-the-art models.
In 13 of the 18 experiments AB training maintains competitive performance compared to the traditional baseline using the same hyperparameters, Figures 3 and 6. Without the independently trained groups, the performance is noticeably reduced, see Figure 3(a), while the compression values are similar, Figure 5.
[…]
Across all experiments, we observed favorable compression ratios (Figures 7 and 5). For runs which reached the same or better loss value, the Vision Transformer’s compression ratios range from 1.89:1 to 2.54:1. ResNet-50’s compression was less impressive, although still respectable (ranging from 1.19:1 to 1.72:1 for outperforming runs). This suggests a potential interaction between model architecture and the effectiveness of our method.
Table 2 demonstrates the effectiveness of AB training compared to existing low-rank and pruning methods. We present both a standard benchmark (ResNet-50 on ImageNet-2012) and an idealized scenario (VGG16 on CIFAR-10). The latter highlights the potential of compression techniques due to VGG16’s over-parameterization and CIFAR-10’s relative simplicity (Chang et al., 2023). AB training outperformed expectations in this setting, achieving a 44.14:1 compression ratio (reducing the model to 2.27% of its original size) with negligible accuracy loss. This surpasses the compression achieved by ICP (3.67% of original size) and ABCPrune (11.26%).
To compare the communication required for training the other model compression methods shown in Table 2 with traditional DP, we estimate their maximum communication savings achievable. As all of these methods iteratively reduce the network size, we assume that the all start with the full-rank representation and remove parameters during training. In out experiments, we found that once the network began to compress the network, it removed much of the network quickly, then removed small amounts during the remaining training steps. However, we do not know the rate at which parameters are removed. Taking this into account, we assume that models train close to full rank for 25% of training then close to their final model state for 75% of training. We describe the estimated communication reduction (ECR) as
(4) $\text{ECR}=100\%-F-Lc$
where $F$ and $L$ are the percentages of training spent near full-rank and the most-compressed network state respectively and $c$ is the final compression ratio expressed as a percentage.
AB training demonstrates the greatest communication reduction in the ResNet-50 benchmark with the ECR. In practice, we measured an approximate reduction of 50% for all use-cases. This discrepancy is caused by other communications needed during training outside of the model synchronization. With the measurement, we outperform all models except OIALR and DLRT, which reduce the number of trainable parameters by training the $\Sigma$ matrix from a network’s weight’s SVD. In the VGG16 scenario, ICP (Chang et al., 2023) and AB training without using individual groups show slightly higher extimated communication reductions than AB training itself. This suggests that AB training offers superior communication efficiency until models reach extremely high compression levels.
While promising, our method also reveals challenges. As expected, we observe generalization improvements in all experiments, these are most noticable in Figure 4. This is likely due to the low-rank representation in combination with the independently trained groups; both of which have regularizing effects (see ‘No Groups’ measurements in Figure 6). However, at larger scales, both loss and accuracy degrade significantly. Combined with the decreasing compression ratios seen with increasing global batch size (Figure 5), this suggests that the averaging of independently trained subgroups becomes detrimental at extreme scales.
Similar trends in the fixed global batch size scenario (Figure 6) suggest an incompatibility between the very large global batch size and the much smaller batches used during the independent training phases. This likely leads to overly large updates in the independent groups, causing them to diverge from each other. Although this divergence increases the regularization effect of independent group training, is reduces the accuracy of the network. Our compression results (Figure 7) support this hypothesis. As the regularization effect strengthens with scale, the divergences between independent groups occur primarily in the smaller, yet important, singular values. When these divergences occur, averaging them together will cause the divergent singular values to shrink relative to the other values.
These challenges likely stem from suboptimal model merging or overly aggressive learning rate schedules. Since it is common practice to adjust the learning rate based on global batch size in traditional DP training, a similar strategy may be required for our method.
Our findings highlight the need for further exploration across diverse models, datasets, and hardware configurations. Understanding the interplay between network bandwidth limitations, and our method’s potential benefits in real-world environments is crucial. Finally, the observed accuracy degradation at scale motivates research into improved update mechanisms, potentially exploring non-average update rules, mixing matrices, or loss-based weighted averaging schemes.
Table 2. Comparison of low-rank and pruning methods for ResNet-50 (on ImageNet-2012) and VGG16 (on CIFAR10). ’Difference to Baseline’ indicates validation top 1 performance relative to the original full-rank model in each study, with positive values denoting better performance. AB training used a global batch size of 4,096 for ImageNet and 1,024 for CIFAR10, achieving maximum top 1 accuracies of 75.67% and 91.87% respectively. The estimated communication reduction (ECR) is defined by Equation 4. AB training’s ECR assumes independent groups do not utilize the compute system’s interconnect. OIALR’s and DLRT’s ECR uses the compression of the trainable parameters as they report.
Method
Difference
to Baseline
Compression
Ratio
ECR
ResNet-50 ImageNet-2012 AB +1.55 % $1.39:1$ 73.29 %*
AB (no groups) -0.02 % $1.27:1$ 15.94 %
OIALR (Coquelin et al., 2024) -1.72 % $1.21:1$ 63.64 %
DLRT (Schotthöfer et al., 2022) -0.56 % $1.85:1$ 64.35 %
PP-1 (Singh et al., 2019) -0.20 % $2.26:1$ 41.81 %
CP (He et al., 2017) -1.40 % $2.00:1$ 37.5 %
SFP (He et al., 2018) -0.20 % $2.39:1$ 43.62 %
ThiNet (Luo et al., 2017) -1.50 % 2.71 : 1 47.32 %
ABCPrune (Lin et al., 2020a) -2.15 % 1.84 : 1 34.24 %
VGG16 CIFAR-10 AB -0.23 % 44.14 : 1 65.60 %*
AB (no groups) -0.89 % 36.63 : 1 72.95 %
OIALR (Coquelin et al., 2024) +0.10 % 3.70 : 1 64.59 %
DLRT (Schotthöfer et al., 2022) -1.89 % 1.79 : 1 16.88 %
ABCPrune (Lin et al., 2020a) +0.06 % 8.83 : 1 66.51 %
ICP (Chang et al., 2023) -0.31 % 27.25 : 1 72.25 %
6. Conclusion
Our experimental results demonstrate the significant potential of our AB training method to reduce communication overheads by utilizing low-rank representations and independently trained worker subgroups. The consistent 50% reduction in network traffic has the potential to unlock new possibilities within distributed-memory computing environments. This could be particularly impactful for scientific research on machines without high-speed interconnects and for exploring even larger, more complex models. Additionally, the pronounced regularization effects observed at smaller scales offer a promising direction for improving generalization and performance.
While our method successfully addresses communication challenges, its performance at extreme scales highlights the complex interplay between low-rank representations, large batch effects, and hyperparameter optimization. Our findings emphasize the need for further research into tailored hyperparameter strategies and novel update mechanisms to fully harness the potential of this approach in massively scaled scenarios.
This work represents a significant step towards communication-efficient distributed training. Our results offer valuable insights and highlight promising areas for future investigation. By addressing the limitations uncovered in our study, we believe the research community can pave the way for even more efficient and scalable training of large-scale neural networks, ultimately furthering scientific discovery across a variety of domains.
Acknowledgements.
This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. This work is supported by the Helmholtz Association Initiative and Networking Fund under the Helmholtz AI platform grant and the HAICORE@KIT partition.

### v2

@@
 AB-Training: A Communication-Efficient Approach for Distributed Low-Rank Learning
-Abstract.
-Communication bottlenecks hinder the scalability of distributed neural network training, particularly on distributed-memory computing clusters.
-To significantly reduce this communication overhead, we introduce AB-training, a novel data-parallel training method that decomposes weight matrices into low-rank representations and utilizes independent group-based training.
-This approach consistently reduces network traffic by 50% across multiple scaling scenarios, increasing the training potential on communication-constrained systems.
-Our method exhibits regularization effects at smaller scales, leading to improved generalization for models like VGG16, while achieving a remarkable 44.14 : 1 compression ratio during training on CIFAR-10 and maintaining competitive accuracy.
-Albeit promising, our experiments reveal that large batch effects remain a challenge even in low-rank training regimes.
-Distributed training, Low-rank representation, Neural networks, Communication efficiency
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 53rd International Conference on Parallel Processing; August 12–15, 2024; Gotland, Swedenccs: Computing methodologies Neural networksccs: Computing methodologies Machine learning algorithmsccs: Networks Network performance analysisccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Batch learning
-Figure 1.
-Average total interconnect traffic measured during the training of a Vision Transformer (ViT) B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ with traditional, synchronous, full-rank data parallel training as implemented by PyTorch’s DistributedDataParallel (DDP) class and our AB training method.
-The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
-[…]
-Average total interconnect traffic measured during the training of a Vision Transformer B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$.
-The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
-1. Introduction
+Abstract
+Communication bottlenecks severely hinder the scalability of distributed neural network training, particularly in high-performance computing (HPC) environments.
+We introduce AB-training, a novel data-parallel method that leverages low-rank representations and independent training groups to significantly reduce communication overhead.
+Our experiments demonstrate an average reduction in network traffic of approximately 70.31% across various scaling scenarios, increasing the training potential of communication-constrained systems and accelerating convergence at scale.
+AB-training also exhibits a pronounced regularization effect at smaller scales, leading to improved generalization while maintaining or even reducing training time.
+We achieve a remarkable 44.14 : 1 compression ratio on VGG16 trained on CIFAR-10 with minimal accuracy loss, and outperform traditional data parallel training by 1.55% on ResNet-50 trained on ImageNet-2012.
+While AB-training is promising, our findings also reveal that large batch effects persist even in low-rank regimes, underscoring the need for further research into optimized update mechanisms for massively distributed training.
+Index Terms:
+neural networks, data parallel training, low-rank decomposition, singular value decomposition
+I Introduction
 The relentless pursuit of predictive performance in machine learning has led to neural networks of unprecedented scale and complexity.
@@
 However, this performance comes at a steep computational cost, as training large models often demands large datasets.
-These datasets often exceed the capabilities of individual devices, necessitating distributed training approaches.
-Data parallelism (DP) is the most common distributed training technique for neural networks (Ben-Nun and Hoefler, 2019).
+The sizes of modern models and datasets necessitates the usage of distributed training approaches.
+Data parallelism (DP) is the most common distributed training technique for neural networks [1].
 It replicates the model across multiple devices, each of which processes a subset of the data, known as the local batch, before aggregating the gradients and updating model parameters.
-A persistent challenge in data parallel (DP) training is the communication bottleneck created by synchronizing large model representations across compute nodes (Niu et al., 2011).
+A persistent challenge in data parallel (DP) training is the communication bottleneck created by synchronizing large model representations across compute nodes [2].
+As the accelerators of neural networks increase in computing power and on-board memory, models finish the forward-backward pass faster.
+In the standard implementations of DP training, the model gradients are eagerly synchronized once they are calculated during the backward pass.
+As models get larger and accelerators get faster, this synchronization happens more frequently and with more data.
+To solve this, one must either increase the available interconnect bandwidth or reduce the rate at which packets are sent to other nodes.
 This bottleneck severely limits scalability, particularly on distributed-memory computing systems with potential network constraints.
-Large batch effects (Goyal et al., 2018), pose an additional challenge for maintaining predictive performance, as accuracy degrades once the global batch size crosses a critical threshold.
-To address these challenges, we propose a novel low-rank training method based on the singular value decomposition (SVD) and a hierarchical training scheme designed to significantly reduce communication overheads in distributed settings.
-Our method leverages the observation that low-rank representations can promote regularization (Coquelin et al., 2024) and minimize the amount of data transferred.
+Furthermore, DP training increases the batch size used to calculate the gradients themselves which, at large scales, reduces the generalizablity of the network [3].
+These large batch effects pose an additional challenge for maintaining predictive performance, as accuracy degrades once the global batch size crosses a critical threshold.
+To address these challenges, we propose a novel low-rank training method based on the singular value decomposition (SVD) and hierarchical training specifically designed to reduce the interconnect traffic in distributed settings.
+Our method leverages the fact that low-rank representations can reduce the amount of data transferred [4], as well as the observation that low-rank representations can promote regularization [5].
 Key contributions of our work include:
-  • We find a 50% reduction in network traffic during distributed training on average as compared to traditional synchronous DP training, enabling experimentation with larger batch sizes and potentially accelerating training within distributed computing environments without high-speed interconnects.
+  • Through the usage of independent training groups and low-rank representations, we reduce the network traffic during distributed training by an average of70.31% as compared to traditional synchronous DP training.
   • Our method demonstrates improved regularization effects at smaller scales, particularly on the Vision Transformer (ViT), leading to better predictive performance on unseen data.
-However, it also sheds light onto the challenge of reducing large batch effects without hyperparameter tuning.
-  • In an ideal scenario, AB training achieves a compression ratio of 44.14 : 1 for the VGG16 (Liu and Deng, 2015) model when trained on CIFAR-10 (Krizhevsky, 2009) while maintaining a competitive accuracy.
+  • In an ideal scenario, AB training achieves a compression ratio of 44.14 : 1 for the VGG16 [6] model when trained on CIFAR-10 [7] while maintaining a competitive accuracy.
 In more standard scenarios, it achieved compression values ranging from 1.19 : 1 to 2.54 : 1 when exceeding the accuracy of traditional DP training.
-Our work highlights promising directions for how to fully harness the potential of low-rank training in massively scaled scenarios, with implications for both distributed-memory computing and the broader ML community.
-2. Related Work
-2.1. Distributed Training of Neural Networks
+Our work highlights promising directions for how to harness the potential of low-rank training in large-scale scenarios, with implications for both distributed-memory computing and the broader ML community.
+We provide an open-source implementation of our method, as well as the configurations to reproduce the experiments shown [^1] Will be made available upon publication.
+II Related Work
+II-A Distributed Training of Neural Networks
 The computational demands of training large neural networks have driven the development of distributed training paradigms, with data parallelism as a cornerstone approach.
-In a traditional data parallel (DP) training scheme, the model is replicated across multiple devices, each processing a subset of the training data independently.
+In a traditional data parallel (DP) training scheme, the model is replicated across multiple devices, each processing a unique subset of the training data independently.
 The data subset that one model instance processes is the local batch, while the global batch is the set of all local batches.
 After each forward-backward pass, gradients are aggregated across all model instances before the model is updated by the optimizer.
-As the number of compute resources and the model size increase, communication overheads incurred during gradient synchronization become a significant constraint, particularly in systems without high-bandwidth interconnects.
-Other parallelization strategies, such as model parallelism, introduce alternative ways to distribute training across devices, but come with their own communication and load-balancing complexities.
-Several techniques have been proposed to mitigate the communication bottleneck in data parallel training, including gradient accumulation, topology-aware communication patterns, asynchronous methods, and gradient compression (Vogels et al., 2019; Abrahamyan et al., 2022).
-Methods like H-SGD (Lin et al., 2020b) and DASO (Coquelin et al., 2022) utilize localized synchronization within smaller groups between global updates, potentially reducing communication costs by leveraging the network’s topology.
-Asynchronous approaches typically make use of a parameter server for aggregating gradients (Niu et al., 2011) and often require extensive hyperparameter tuning.
+As the number of compute resources and the model size increase, communication overheads incurred during gradient synchronization become a significant constraint.
+Other parallelization strategies, such as model parallelism, introduce alternative ways to distribute training across devices, but come with their own communication and load-balancing challenges.
+Several techniques have been proposed to mitigate the communication bottleneck in data parallel training, including gradient accumulation, topology-aware communication patterns, asynchronous methods, and gradient compression [4, 8].
+Methods like H-SGD [9] and DASO [10] utilize localized synchronization within smaller groups between global updates, reducing communication costs by leveraging the system’s network topology.
+Asynchronous approaches typically make use of a parameter server for aggregating gradients [2] and often require extensive hyperparameter tuning [1].
 Despite these advancements, the scalability of data parallelism remains constrained by communication bottlenecks and large batch effects, where increasing the global batch size negatively impacts generalization performance.
-2.2. Low-Rank Neural Network Training
+II-B Low-Rank Neural Network Training
 A promising avenue for reducing both the memory footprint and computational cost of neural networks lies in low-rank training.
 Singular Value Decomposition (SVD) plays a crucial role in this domain.
-By factoring a weight matrix $W$ into three matrices, $U$, $\Sigma$, and $V^{T}$, and retaining only the $k$ largest singular values, a low-rank approximation with significantly fewer parameters can be obtained given a sufficiently large $k$ value.
-Note that aggressive compression can lead to a loss of approximation quality.
+By factoring a neural network’s weight matrix $W$ into three matrices (an orthogonal basis $U$, a diagonal matrix of singular values $\Sigma$, and an orthogonal basis $V^{T}$) and retaining only the $k$ largest singular values, a low-rank approximation with significantly fewer parameters can be obtained given a sufficiently large $k$.
+However, aggressive compression can lead to a loss of approximation quality.
 Several methods leverage SVD for compressing neural networks during training.
-Some train directly on the $U$, $\Sigma$, and $V$ matrices (Schotthöfer et al., 2022; Coquelin et al., 2024), while others create different representations using the SVD and training on them (Liebenwein et al., 2021; Yang et al., 2020a).
-Both approaches can offer regularization-like benefits, potentially improving generalization (Winata et al., 2020; Cahyawijaya et al., 2021; Phan et al., 2020).
-Gradient compression using SVD has found application in federated learning scenarios (Wang et al., 2023b), where bandwidth constraints necessitate reduced communication.
+Some train directly on the $U$, $\Sigma$, and $V$ matrices [11, 5], while others create different representations using the SVD and training on them [12, 13].
+Both approaches can offer regularization-like benefits, potentially improving generalization [14, 15, 16].
+LoRA [17], and methods similar to and inspired by it, utilize low-rank representations alongside the frozen full-rank weights for fine-tuning.
+Gradient compression using SVD has found application in federated learning scenarios [18], where bandwidth constraints necessitate reduced communication.
 While these techniques address computational challenges in neural network training, their integration within distributed environments to specifically optimize communication remains an active area of research.
-Low-rank factorizations can heavily reduce the communication required during distributed training.
-Existing approaches simply train low-rank models in parallel, but do not make any efforts to use the distributed system to their advantage (Coquelin et al., 2024; Wang et al., 2023a, 2021).
-To the best of our knowledge, there exist no methods that leverage both distributed architectures and low-rank representations to train neural networks.
-3. Methodology
+More precisely, existing approaches simply train low-rank models in parallel, but do not make any efforts to use the distributed system to their advantage [5, 19, 20].
+II-C Pruning
+Pruning aims to reduce model size by removing irrelevant parameters.
+It is the primary competition to low-rank methods for reducing the size of neural networks.
+One well-known method is Iterative Magnitude Pruning (IMP), which involves repeatedly training a network, removing small-magnitude weights, and retraining from the initial state[21].
+While this process can result in highly compressed models., unstructured pruning methods like IMP often lead to irregular sparsity patterns, making it difficult to leverage specialized hardware optimized for sparse computation.
+Structured pruning techniques target specific components of the network.
+For example, in convolutional layers removing filters [22, 23, 24] or channels [25, 26, 27] can maintain traditional dense parameter layouts to make use of efficient, pre-existing acceleration methods.
+III Methodology
 Data parallel training of neural networks is constrained by the the communication bottleneck caused by synchronizing massive model representations across compute nodes and large batch effects, a phenomenon where accuracy degrades at large global batch sizes.
-In an attempt to tackle these challenges, our proposed training method integrates low-rank weight representations within a data parallel framework.
-It leverages the reduced network capacity stemming from topologically informed communication patterns and low-rank training while escaping local minima using group training.
-We utilize SVD to decompose weight matrices into low-rank representations, significantly reducing the amount of data communicated during training. We decompose the weights of a given layer as
-(1) $W_{m\times n}=A_{m\times k}B_{k\times n}=\left(U_{m\times k}\Sigma_{k\times k}^{1/2}\right)\left(\Sigma_{k\times k}^{1/2}V_{k\times n}^{\top}\right)$
-where $U$, $\Sigma$, and $V$ are determined by the SVD of $W$ and $k<\texttt{min}\left(m,n\right)$.
-Using a low-rank representation in training has the added benefit of reducing overfitting, a common problem in training neural networks (Yang et al., 2020b; Hu et al., 2021).
-Large batch effects can be trickier to deal with. Consider the standard batch stochastic gradient descent (SGD) optimization step
-(2) $W_{t}=W_{t-1}-\frac{\eta}{N}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{t-1}\right),$
+In an attempt to tackle these challenges, our proposed training method integrates low-rank weight representations within a hierarchical data parallel framework.
+It leverages the reduced network capacity stemming from topologically informed communication patterns and low-rank training while also escaping local minima using group training.
+Working with large batches can have unintended consequences. Consider the standard batch stochastic gradient descent (SGD) optimization step
+$W_{t}=W_{t-1}-\frac{\eta}{B}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{t-1}\right),$ (1)
 where $W_{t}$ is a parameter at time $t$ in training, $\eta$ is the learning rate, $N$ is the batch size, and $Q_{i}\left(W\right)$ is the value of the loss function for the $i$-th data element.
 It is easily shown that for a given iteration $k>0$, the weights are simply the sum of all the previous update steps or
-(3) $W_{k}=W_{0}-\frac{\eta}{B}\sum_{s=1}^{k-1}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{s}\right)$
-With a larger batch size, $B$, the magnitude of the gradients shrinks.
-This removes the randomness inherent to SGD’s success and makes the network more prone to getting stuck in local minima (Chaudhari et al., 2019).
-To mitigate this effect, we draw motivation from ensemble approaches.
-By dividing the DP model replicas into subgroups to be trained on independent data subsets, we essentially form an ensemble of models.
-Individually, these groups converge towards different local minima compared to full-system training, promoting exploration of the loss landscape.
-After training in subgroups, we average the independently trained models.
-This merging strategy has seen some success in federated learning methods (Giuseppi et al., 2022).
-Compared to federated use-cases, which typically deal with models trained independently on biased datasets, AB training benefits from randomly sampling from the same dataset and thus has unbiased batches during our independent training phases.
-During the independent training phases, each group trains on either $A$ or $B$, while the other component remains fixed.
-As the orthogonal component of $W$ is stable later in training (Coquelin et al., 2024), we only train on one of $A$ or $B$ in an attempt to reduce the divergence between the independent groups.
-At the end of the independent training phase, we average $A$ and $B$ across all workers to merge the independently trained models together.
+$W_{k}=W_{0}-\frac{\eta}{B}\sum_{s=1}^{k-1}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{s}\right)$ (2)
+With a larger batch size, $B$, the magnitude of the gradients shrinks. This removes the randomness inherent to SGD’s success and makes the network more prone to getting stuck in local minima [28, 29].
+To mitigate this effect, we draw motivation from the fact that the noise inherent to the gradients in minibatch SGD is an important factor in why it generalizes better than large batch gradient descent [29].
+During training we divide the workers into small sub-groups; each of which train the model independently for a number of steps.
+By dividing the DP model replicas into subgroups to be trained on independent data subsets, we effectively form an ensemble of models.
+Individually, these groups benefit from the improved generalization of minibatch SGD by using the noisier gradients to explore of the loss landscape to find minima which would otherwise be missed by large-batch SGD.
+Additionally, training in independent groups heavily reduces the network traffic.
+To further reduce the amount of data communicated during training, we utilize SVD to decompose weight matrices into low-rank representations. We represent the weights of a given layer as
+$W_{m\times n}=A_{m\times k}B_{k\times n}=\left(U_{m\times k}\Sigma_{k\times k}^{1/2}\right)\left(\Sigma_{k\times k}^{1/2}V_{k\times n}^{\top}\right)$ (3)
+where $U$, $\Sigma$, and $V$ are determined by the SVD of $W$ and $k<\texttt{min}\left(m,n\right)$.
+In the case that the weight matrix has more than two dimensions, the trailing dimensions are flattened to form a two-dimensional representation to be used as $W$.
+If the shape of the matrix $W$ is unfavorable ($n>m$) we transpose $W$ before the decomposition.
+One of the most obvious methods for training a low-rank network would be to train directly on the $\Sigma$ matrix.
+However, it is not beneficial to have any trainable weights set to zero during training as these connections are effectively dead during backpropagation.
+Since the off-diagonal values of $\Sigma$ (as dictated by SVD) are zero, training on only $\Sigma$ will have the effect of freezing the orthogonal vectors of $U$ and $V$.
+To allow the network to modify the orthogonal vectors and to maintain a dense representation, we decompose the network as shown in Equation 3.
+Interestingly, training a network using low-rank representations has the added benefit of reducing overfitting, a common problem where the performance on the training dataset is much higher than on the validation or test datasets [13, 30].
+During the independent training phase, half of the sub-groups train $A$ while the others train $B$; in this phase the other component remains fixed.
+This ensures that at least a portion of the orthogonal vectors inherent to $W$ are fixed in time, reducing the chance that the groups greatly diverge and increases the gradient noise due to the smaller batch size.
+With this method, the groups can more easily learn to modify the full low-rank representations of the network’s parameters.
+At the end of the independent training phase, we average $A$ and $B$ across all workers to merge the independently trained models.
+This merging strategy has seen success in federated learning [31], a use-case which typically deals with models trained independently on biased datasets.
+As we use unbiased batches during the independent training phase, we expect the negative aspects of this average to be minimal.
 Our AB training procedure consists of the following phases:
-Distributed warmup:
-Initial training with full-rank data parallelism allows the weights to adapt rapidly away from their random initialization.
+  • Full Rank DP Warmup: Initial training with full-rank data parallelism allows the weights to move rapidly away from their random initialization.
 To avoid early divergence, we employ a traditional, full-rank data parallel warmup phase for a given number of iterations (warmupSteps).
-Independent AB decomposition:
-Each model instance independently computes an AB decomposition of its weight matrices as per Equation 1.
+  • Independent AB decomposition: Each model instance independently computes an AB decomposition of its weight matrices as per Equation 3.
 This AB decomposition allows us to represent the weight matrices more compactly, reducing the amount of data to be communicated during training.
-The rank of this approximation is determined by a user-provided hyperparameter (sigmaCutoff) times the largest singular value.
-Values smaller than this are removed.
-Group-only training:
-Half of the independent groups will train the $A$ matrix, denoted as the $A$ groups, while the other groups will train $B$, the $B$ groups.
-For a given number of iterations (numABSteps), processes in the $A$ group only train the $A$ matrices, while processes in the $B$ group only train the $B$ matrices, the matrix not being trained remains frozen in this phase.
-Synchronization and update:
-We average both the trained and frozen $A$ and $B$ matrices independently to synchronize the model instances across all processes.
-This is crucial for ensuring that the overall model remains consistent across all compute nodes.
+The rank of this approximation is determined by a user-provided hyperparameter (sigmaCutoff) and the largest singular value.
+Values smaller than the largest singular value times sigmaCutoff are removed.
+  • Group training: Half of the independent groups will train the $A$ matrix, denoted as the $A$ groups, while the other groups will train $B$, the $B$ groups.
+For a given number of iterations (numABSteps), the $A$ group only trains the $A$ matrices and the $B$ group only trains the $B$ matrices, the matrices not being trained are frozen.
+  • Synchronization and update: We average both the trained and frozen $A$ and $B$ matrices to synchronize the model instances across all processes.
 Afterwards, we reconstruct the full-rank weight matrices from the updated $A$ and $B$ components.
-Full-rank rebound:
-We then train the full-rank network with traditional data parallelism for a short period, which we refer to as the fullRankReboundSteps.
+  • Full-rank rebound: We then train the full-rank network with traditional data parallelism for a number of iterations (fullRankReboundSteps).
 This phase promotes convergence and helps mitigate potential accuracy degradation observed in low-rank training scenarios.
-A formalized algorithm is shown in Algorithm 1 and the method diagram is shown in Figure 2.
-Figure 2. The AB training procedure.
-[…]
-The AB training procedure. See pseudocode or description in text.
+A formalized algorithm is shown in Algorithm 1 and a method diagram is shown in Figure 1.
+Figure 1: A UML diagram of the AB training procedure.
 Input : Model, $M$ , training data, $numTrainingSteps$, hyperparameters ($warmupSteps$, $numABSteps$, $fullRankReboundSteps$)
@@
 10       foreach $W$ in $M$ do
-11             $A,B\leftarrow\texttt{abDecomposition}(W)$triangleright Equation 1
+11             $A,B\leftarrow\texttt{abDecomposition}(W)$triangleright Eq. 3
 12             removeSmallSingularValues()
@@
 Algorithm 1 The AB training method. $W$ is a parameter of the input model $M$ and $worldSize$ is the number of workers used for traditional DP training.
-Popular neural network optimizers like AdamW (L.
-and H., 2018) rely on second-order derivative approximations.
-When we change the network’s structure (e.g., between low-rank and full-rank representations), these optimizer states become incompatible due to their dimensionally.
-To address this, we use a “learning rate rebound” strategy.
+Popular neural network optimizers like AdamW [32] rely on second-order derivative approximations.
+When we change the network’s structure, e.g., between low-rank and full-rank representations, the optimizer states used in these approximations become incompatible due to their dimensionally.
+To address this, we use a learning rate rebound strategy.
 This involves reducing the learning rate to near-zero and gradually increasing it back to its scheduled value over a user-provided number of steps, allowing the optimizer to adjust to the new parameter representations.
-We do a learning rate rebound after each time the network’s parameters change shape.
-4. Experimental Evaluation
-To evaluate the effectiveness of our AB training approach, we conducted experiments on established neural network architectures.
-Our primary objectives were to demonstrate significant reductions in communication overhead, potential regularization benefits, scalability, and achievable compression during training.
-We focused on the ResNet-50 (He et al., 2016) and Vision Transformer (Dosovitskiy et al., 2021) B/16 models trained on the ImageNet-2012 dataset (Krizhevsky et al., 2012) for image classification as well as VGG16 (Liu and Deng, 2015) trained on CIFAR-10 (Krizhevsky, 2009).
-We compare our results against traditional DP training as implemented by PyTorch’s DDP package, referred to as ‘Traditional DDP’ or ‘Trad.
+We do a learning rate rebound whenever the network’s parameters change shape.
+IV Experimental Evaluation
+To evaluate the effectiveness of our AB training approach, we conducted experiments on well-studied neural network architectures.
+Our primary objectives were to demonstrate significant reductions in interconnect traffic, potential regularization benefits, scalability, and achievable compression during training.
+We focused on the ResNet-50 [33] and Vision Transformer [34] B/16 models trained on the ImageNet-2012 dataset [35] for image classification as well as VGG16 [6] trained on CIFAR-10 [7].
+We compare our results against traditional full-rank DP training as implemented by PyTorch’s DistributedDataParallel (DDP) class, referred to as ‘Traditional DDP’ or ‘Trad.
 DDP’, using the same hyperparameters (HPs).
-We also show results for the AB training setup without any individual trained groups which are denoted as ‘No Group’.
-In these experiments, the $B$ matrix is trained using traditional DP training in place of the full-rank weights.
-The ViT HPs are those used in the original source, while the ResNet-50 parameters utilize the same learning rate scheduler and the default values for the AdamW (L.
-and H., 2018) optimizer.
-The VGG16 HPs are as shown in (Coquelin et al., 2024).
+In these experiments, the $B$ matrix is trained during the group phase using traditional DP training.
+The ViT HPs are those used in the original source, while the ResNet-50 parameters utilize the same learning rate scheduler as the ViT and the default values for the AdamW [32] optimizer.
+The VGG16 HPs are as shown in [5].
 We analyze our training method using two scaling strategies: constant local batch size scaling and constant global batch size scaling.
@@
 In constant global batch size scaling, we maintain a constant global batch size while proportionally decreasing the local batch size as we increase the number of GPUs.
-This emphasizes the communication overhead reductions achieved by our method.
+This emphasizes the interconnect traffic reduction achieved by our method.
+To disentangle the effects of training with independent groups and those of low-rank training, we show measurements for AB training with and without groups.
+For the no-group measurements, the workers train the $B$ matrix with traditional DP methods and denoted as ‘AB - No Groups’.
+At a fixed scale, we compare AB training with and without groups to other pruning and low-rank methods.
 Measurements represent the average of three runs, each of which utilizes a different random seed.
-Models are initialized using orthogonal initialization (Hu et al., 2020).
-4.1. Computational Environment
+Models are initialized using orthogonal initialization [36].
+IV-A Computational Environment
 […]
-4.2. Datasets and Models
+IV-B Datasets and Models
 We used the ImageNet-2012 dataset, which contains 1.2 million images, for our scaling experiments.
 Basic image augmentation was applied, including normalization, random resizing, random cropping, and a random flip.
-We trained two models on this dataset: ResNet-50 (He et al., 2016) and Vision Transformer B/16 (ViT) (Dosovitskiy et al., 2021), chosen for their widespread use and distinct architectures.
-We also trained VGG16 on the CIFAR-10 dataset, which contains 50,000 images. Basic image augmentation was applied in the same way as for the ImageNet experiments.
-All models were trained with the AdamW (Loshchilov and Hutter, 2019) optimizer. Other training hyperparameters can be found in Appendix 1.
-4.3. Hyperparameter Considerations
-Careful hyperparameter tuning is essential for balancing the communication reduction achieved through our method with potential impacts on training stability and accuracy.
-Key hyperparameters (HPs) include the warmup and full-rank rebound phase durations, the number of AB training iterations, the AB decomposition rank-reduction parameter, and the frequency of SVD and synchronization steps.
+We trained two models on this dataset: ResNet-50 [33] and Vision Transformer B/16 (ViT) [34], chosen for their widespread use and distinct architectures.
+We also trained VGG16 on the CIFAR-10 dataset.
+Basic image augmentation was applied in the same way as for the ImageNet experiments.
+All models were trained with the AdamW [38] optimizer.
+Other training hyperparameters can be found in Appendix 1.
+IV-C Hyperparameter Considerations
+As with any method, hyperparameter tuning is essential to performance.
+Key hyperparameters (HPs) of AB training include the warmup and full-rank rebound phase durations, the number of AB training iterations, the AB decomposition rank-reduction parameter, and the frequency of SVD and synchronization steps.
 Additionally, the latter will influence the trade-offs between regularization and communication efficiency.
-Our hyperparameter search on CIFAR-100 (Krizhevsky, 2009) with a minified Vision Transformer model (patch size of eight instead of 16, six heads instead of twelve, depth of six instead of twelve) using Propulate (Taubert et al., 2023) suggested the following guidelines:
+A hyperparameter search on CIFAR-100 [7] with a minified Vision Transformer model (patch size of eight instead of 16, six heads instead of twelve, depth of six instead of twelve) was conducted using propulate [39] to determine the ideal settings for AB-Training’s hyperparameters.
+Using Weights and Biases [40], we measures the importance of these parameters via a Bayesian analysis of the completed runs.
+We find that the most important of these was the numABStesps.
+Controlling how many steps to train in low rank was crucial.
+However, the correlation of this variable to the final top-1 accuracy was only 0.225, indicating that there is a sweet spot between too many and too few steps.
+The number of warmup steps was also found to be quite important, with short warmups performing worse.
+The search and our analysis of the results suggested the following guidelines which are used for all experiments:
   • warmupSteps: 20% of the total training steps
@@
   • Learning rate rebound steps: $0.5\cdot\texttt{numABSteps}$
-4.4. Constant Local Batch Size Scaling
+IV-D Constant Local Batch Size Scaling
 In this experiment, we maintain a fixed local batch size while increasing the available compute resources and decreasing the number of training steps per epoch, resulting in an increasing global batch size but less training iterations.
@@
 If this is the case, then AB training without independent groups will begin to outperform AB training with groups.
-Figure 3 shows the highest top 1 accuracy achieved by each model during training.
-Figure 4 shows the lowest binary cross-entropy loss values achieved across all epochs.
-The compression values for the trained models are shown in Figure 5.
+Figure 2 shows the highest top 1 accuracy achieved by each model during training. The compression values for the trained models are shown in Figure 4(a).
 (a) Vision Transformer B/16
 (b) ResNet-50
-Figure 3.
-Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256.
+Figure 2: Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256.
 Global batch sizes range from 2,048 to 32,768 in powers of 2.
 Error are plotted, though not always visible.
-[…]
-Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256.
-Global batch sizes range from 2,048 to 32,768 in powers of 2.
+IV-E Constant Global Batch Size Scaling
+In this experiment, we maintain a fixed global batch size while scaling compute resources.
+This contrasts with the standard approach of keeping local batch sizes constant, emphasizing the interconnect traffic reductions achieved by AB training.
+This will reduce the number of computations per process per forward-backward pass.
+With a fixed global batch size of 4,096, a four-node run would have a local batch size of 256, whereas an eight-node run would have a local batch size of 128.
+This experiment serves to show the performance of AB training as the time required to process a local batch shrinks, as is the case when the accelerators on a system are replaced but the network infrastructure remains the same.
+It also provides insight into the ability of independent training groups to learn meaningful representations when the batch size used for independent training is much smaller than the global batch size.
+Without the independent groups, the accuracy of the trained network should be similar throughout scaling.
+Results for these runs are shown in Figures 3, 4(b) and 5.
 (a) Vision Transformer B/16
 (b) ResNet-50
-Figure 4.
-Lowest binary cross-entropy (the loss function used during training) measurements during each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256.
-Global batch sizes range from 2,048 to 32,768 in powers of 2.
+Figure 3: Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
 Error are plotted, though not always visible.
-[…]
-Lowest binary cross-entropy (the loss function used during training) measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256.
-Global batch sizes range from 2,048 to 32,768 in powers of 2.
-AB training is shown to have a regularization effect for both ResNet-50 and the Vision Transformer.
-Large batch effects are also visible as the loss values get worse (larger) as the global batch size increases.
-Figure 5.
-Average compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
-[…]
-Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
-4.5. Constant Global Batch Size Scaling
-In this experiment, we maintain a fixed global batch size while scaling compute resources.
-This contrasts with the standard approach of keeping local batch sizes constant, emphasizing the communication overhead reductions achieved by AB training.
-With a fixed global batch size of 4,096, a four-node run would have a local batch size of 256, whereas an eight-node run would have a local batch size of 128.
-This experiment directly quantifies the communication benefits of our method in scenarios where local batch sizes become very small due to scaling.
-It also provides insight into the ability of independent training groups to learn meaningful representations when the batch size used for independent training is much smaller than the global batch size.
-Results for these runs are shown in Figures 6 and 7.
-(a) Vision Transformer B/16
-(b) ResNet-50
-Figure 6.
-Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
-Error are plotted, though not always visible.
-[…]
-Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
-Figure 7.
-Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
-[…]
-Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
-5. Discussion
-Table 1.
-Average interconnect traffic, in GB/s, on each processes.
-$N$ represents either the number of nodes for the constant global batch size experiments or the global batch size in thousands for the constant local batch size experiments.
-Constant Local
-Batch Size
-Constant Global
-Batch Size
-$N$ Baseline AB Baseline AB
-ViT 2 $4.67\pm 0.00$ $2.10\pm 0.07$
-4 $4.77\pm 0.06$ $2.12\pm 0.01$ $4.81\pm 0.05$ $1.86\pm 0.39$
-8 $4.58\pm 0.00$ $1.76\pm 0.29$ $7.05\pm 1.03$ $3.45\pm 0.18$
-16 $4.44\pm 0.10$ $1.74\pm 0.07$ $13.66\pm 0.16$ $5.33\pm 0.70$
-32 $4.07\pm 0.20$ $1.58\pm 0.24$ $14.09\pm 3.47$ $7.88\pm 0.02$
-ResNet-50 2 $2.06\pm 0.01$ $1.02\pm 0.00$
-4 $2.07\pm 0.01$ $0.98\pm 0.05$ $2.06\pm 0.01$ $1.00\pm 0.01$
-8 $2.00\pm 0.05$ $1.05\pm 0.05$ $2.67\pm 0.05$ $1.16\pm 0.14$
-16 $1.96\pm 0.12$ $1.15\pm 0.00$ $3.28\pm 0.25$ $1.39\pm 0.21$
-32 $1.62\pm 0.03$ $1.05\pm 0.01$ $3.50\pm 1.00$ $1.76\pm 0.47$
-Across both scaling scenarios, we find a consistent reduction of approximately 50% on average network traffic, demonstrating the effectiveness of our AB training method in reducing communication overheads.
-Comprehensive results are listed in Table 1, which further highlights the significant bandwidth requirements for training even moderately sized neural networks, a challenge that would be exacerbated for larger, state-of-the-art models.
-In 13 of the 18 experiments AB training maintains competitive performance compared to the traditional baseline using the same hyperparameters, Figures 3 and 6.
-Without the independently trained groups, the performance is noticeably reduced, see Figure 3(a), while the compression values are similar, Figure 5.
-[…]
-Across all experiments, we observed favorable compression ratios (Figures 7 and 5).
-For runs which reached the same or better loss value, the Vision Transformer’s compression ratios range from 1.89:1 to 2.54:1.
+(a) Constant local batch size of 256.
+(b) Constant global batch size of 4,096.
+Figure 4: Compression ratios for AB training with and without groups and a traditional DP baseline on ImageNet-2012 for two network architectures
+(a) Scaled interconnect traffic. The ‘Peak Bandwidth’ is that of each port of each node’s Infiniband card.
+(b) Training wall-clock time.
+Figure 5: Scaled interconnect traffic and job wall-clock time for the ViT B/16 trained on ImageNet-2012.
+Scaling is based on the average time required for the calculation and communication of gradients on 4 nodes.
+V Discussion
+Table II shows that our AB training method consistently achieves an average reduction of 70.31% in scaled network traffic across the scaling tests.
+The scaled network traffic is calculated as the average network traffic across the entire job scaled by the percentage of time required by the backward step during training (measured as 53.84% on four nodes).
+The traffic reduction highlights AB training’s effectiveness in mitigating the DP training communication bottleneck, a key concern in distributed training, especially for large models in HPC environments.
+These results underscore the significant bandwidth requirements for training even moderately sized neural networks, a challenge further exacerbated by larger models.
+AB training maintains competitive accuracy compared to traditional DP training in 13 of the 18 scaling experiments (Figures 2 and 3), using the same hyperparameters and in a similar time frame.
+Removing the independently trained groups noticeably reduces accuracy (Figure 2(a)), while compression values remain similar (Figure 4(a)).
+This demonstrates that AB training’s performance benefits stem from both the low-rank representation and the independent subgroup training.
+Our large-scale experiments reveal the interplay between communication efficiency and model accuracy.
+In our ViT experiments with 32 nodes and a constant global batch size (Figure 5), we observe the aforementioned communication bottleneck, where the training time no longer decreases despite additional compute resources.
+However, AB training continues to see a reduction in time to train due to its reduced communication demands, even with the additional overhead stemming from the computation of the SVD for all model weights.
+ResNet-50’s performance plateaus earlier due to its computation becoming the bottleneck, though AB training and the traditional DP baseline have similar accuracy and require roughly the same time to train.
+AB training maintains or reduces the training time as compared to PyTorch’s DDP for all scaling measurements (see Table II).
+This is most obvious in the measurements shown in Figure 5, .
+Across all experiments, we observed favorable compression ratios for AB training(Figures 4(b) and 4(a) and Table II).
+For runs which equaled or exceeded the baseline network accuracy, the Vision Transformer’s compression ratios range from 1.89 : 1 to 2.54 : 1.
 ResNet-50’s compression was less impressive, although still respectable (ranging from 1.19 : 1 to 1.72 : 1 for outperforming runs).
 This suggests a potential interaction between model architecture and the effectiveness of our method.
-Table 2 demonstrates the effectiveness of AB training compared to existing low-rank and pruning methods.
-We present both a standard benchmark (ResNet-50 on ImageNet-2012) and an idealized scenario (VGG16 on CIFAR-10).
-The latter highlights the potential of compression techniques due to VGG16’s over-parameterization and CIFAR-10’s relative simplicity (Chang et al., 2023).
-AB training outperformed expectations in this setting, achieving a 44.14:1 compression ratio (reducing the model to 2.27% of its original size) with negligible accuracy loss.
-This surpasses the compression achieved by ICP (3.67% of original size) and ABCPrune (11.26%).
-To compare the communication required for training the other model compression methods shown in Table 2 with traditional DP, we estimate their maximum communication savings achievable.
-As all of these methods iteratively reduce the network size, we assume that the all start with the full-rank representation and remove parameters during training.
-In out experiments, we found that once the network began to compress the network, it removed much of the network quickly, then removed small amounts during the remaining training steps.
+Table I demonstrates AB training’s effectiveness compared to existing low-rank and pruning methods.
+We showcase two standard benchmarks for these methods: ResNet-50 on ImageNet-2012 and VGG16 on CIFAR-10.
+The latter highlights the potential of compression techniques due to VGG16’s over-parameterization and CIFAR-10’s relative simplicity [27].
+AB training outperformed expectations on CIFAR10, achieving a 44.14 : 1 compression ratio (reducing the model to 2.27% of its original size) with negligible accuracy loss.
+This surpasses the compression achieved by ICP [27] (3.67% of original size) and ABCPrune [25] (11.26%).
+For the ResNet-50 benchmark, AB training was the only one of the listed methods to outperform the baseline.
+To compare the communication required for training the other model compression methods shown in Table I with traditional DP, we estimate their maximum communication savings achievable.
+As all of these methods iteratively reduce the network size, we assume that they all start with the full-rank representation and remove parameters during training.
+In our experiments, we found that once the network began to compress the network, it removed much of the network quickly, then removed small amounts during the remaining training steps.
 However, we do not know the rate at which parameters are removed.
@@
 We describe the estimated communication reduction (ECR) as
-(4) $\text{ECR}=100\%-F-Lc$
+$\text{ECR}=100\%-F-Lc$ (4)
 where $F$ and $L$ are the percentages of training spent near full-rank and the most-compressed network state respectively and $c$ is the final compression ratio expressed as a percentage.
-AB training demonstrates the greatest communication reduction in the ResNet-50 benchmark with the ECR.
-In practice, we measured an approximate reduction of 50% for all use-cases.
-This discrepancy is caused by other communications needed during training outside of the model synchronization.
-With the measurement, we outperform all models except OIALR and DLRT, which reduce the number of trainable parameters by training the $\Sigma$ matrix from a network’s weight’s SVD.
-In the VGG16 scenario, ICP (Chang et al., 2023) and AB training without using individual groups show slightly higher extimated communication reductions than AB training itself.
-This suggests that AB training offers superior communication efficiency until models reach extremely high compression levels.
-While promising, our method also reveals challenges.
-As expected, we observe generalization improvements in all experiments, these are most noticable in Figure 4.
-This is likely due to the low-rank representation in combination with the independently trained groups; both of which have regularizing effects (see ‘No Groups’ measurements in Figure 6).
-However, at larger scales, both loss and accuracy degrade significantly.
-Combined with the decreasing compression ratios seen with increasing global batch size (Figure 5), this suggests that the averaging of independently trained subgroups becomes detrimental at extreme scales.
-Similar trends in the fixed global batch size scenario (Figure 6) suggest an incompatibility between the very large global batch size and the much smaller batches used during the independent training phases.
-This likely leads to overly large updates in the independent groups, causing them to diverge from each other.
-Although this divergence increases the regularization effect of independent group training, is reduces the accuracy of the network.
-Our compression results (Figure 7) support this hypothesis.
-As the regularization effect strengthens with scale, the divergences between independent groups occur primarily in the smaller, yet important, singular values.
-When these divergences occur, averaging them together will cause the divergent singular values to shrink relative to the other values.
-These challenges likely stem from suboptimal model merging or overly aggressive learning rate schedules.
-Since it is common practice to adjust the learning rate based on global batch size in traditional DP training, a similar strategy may be required for our method.
-Our findings highlight the need for further exploration across diverse models, datasets, and hardware configurations.
-Understanding the interplay between network bandwidth limitations, and our method’s potential benefits in real-world environments is crucial.
-Finally, the observed accuracy degradation at scale motivates research into improved update mechanisms, potentially exploring non-average update rules, mixing matrices, or loss-based weighted averaging schemes.
-Table 2.
-Comparison of low-rank and pruning methods for ResNet-50 (on ImageNet-2012) and VGG16 (on CIFAR10).
+For AB training, the percentage of training spent near the final compression state, $L$, is reduced by the percentage time spent in the independent group training phase.
+TABLE I: Comparison of low-rank and pruning methods for ResNet-50 (on ImageNet-2012) and VGG16 (on CIFAR10).
 ’Difference to Baseline’ indicates validation top 1 performance relative to the original full-rank model in each study, with positive values denoting better performance.
@@
 AB training’s ECR assumes independent groups do not utilize the compute system’s interconnect.
-OIALR’s and DLRT’s ECR uses the compression of the trainable parameters as they report.
 Method
@@
 ECR
-ResNet-50 ImageNet-2012 AB +1.55 % $1.39:1$ 73.29 %*
-AB (no groups) -0.02 % $1.27:1$ 15.94 %
-OIALR (Coquelin et al., 2024) -1.72 % $1.21:1$ 63.64 %
-DLRT (Schotthöfer et al., 2022) -0.56 % $1.85:1$ 64.35 %
-PP-1 (Singh et al., 2019) -0.20 % $2.26:1$ 41.81 %
-CP (He et al., 2017) -1.40 % $2.00:1$ 37.5 %
-SFP (He et al., 2018) -0.20 % $2.39:1$ 43.62 %
-ThiNet (Luo et al., 2017) -1.50 % 2.71 : 1 47.32 %
-ABCPrune (Lin et al., 2020a) -2.15 % 1.84 : 1 34.24 %
-VGG16 CIFAR-10 AB -0.23 % 44.14 : 1 65.60 %*
-AB (no groups) -0.89 % 36.63 : 1 72.95 %
-OIALR (Coquelin et al., 2024) +0.10 % 3.70 : 1 64.59 %
-DLRT (Schotthöfer et al., 2022) -1.89 % 1.79 : 1 16.88 %
-ABCPrune (Lin et al., 2020a) +0.06 % 8.83 : 1 66.51 %
-ICP (Chang et al., 2023) -0.31 % 27.25 : 1 72.25 %
-6. Conclusion
-Our experimental results demonstrate the significant potential of our AB training method to reduce communication overheads by utilizing low-rank representations and independently trained worker subgroups.
-The consistent 50% reduction in network traffic has the potential to unlock new possibilities within distributed-memory computing environments.
+ResNet-50 ImageNet-2012 AB +1.55 % $1.39:1$ 73.29 %
+AB - No Groups -0.02 % $1.27:1$ 15.94 %
+OIALR [5] -1.72 % $1.21:1$ 63.64 %
+DLRT [11] -0.56 % $1.85:1$ 64.35 %
+PP-1 [22] -0.20 % $2.26:1$ 41.81 %
+CP [26] -1.40 % $2.00:1$ 37.5 %
+SFP [23] -0.20 % $2.39:1$ 43.62 %
+ThiNet [24] -1.50 % 2.71 : 1 47.32 %
+ABCPrune [25] -2.15 % 1.84 : 1 34.24 %
+VGG16 CIFAR-10 AB -0.23 % 44.14 : 1 65.60 %
+AB - No Groups -0.89 % 36.63 : 1 72.95 %
+OIALR [5] +0.10 % 3.70 : 1 64.59 %
+DLRT [11] -1.89 % 1.79 : 1 16.88 %
+ABCPrune [25] +0.06 % 8.83 : 1 66.51 %
+ICP [27] -0.31 % 27.25 : 1 72.25 %
+AB training demonstrates the greatest communication reduction in the ResNet-50 benchmark with the ECR.
+In practice, we measured an approximate network traffic reduction of 70% for all use-cases, landing quite close to the ECR.
+With the ECR, we outperform all models except OIALR and DLRT, which reduce the number of trainable parameters by training the $\Sigma$ matrix from a network’s weight’s SVD.
+In the VGG16 scenario, ICP [27] and AB training without using individual groups show slightly higher estimated communication reductions than AB training itself.
+This suggests that AB training offers superior communication efficiency until models reach extremely high compression levels.
+Our results also reveal a complex interplay between low-rank representations, large batch sizes, and generalization.
+We observe significant generalization improvement in all experiments, likely due to the regularization effect of low-rank representations and independent group training (see ‘No Groups’ measurements in Figure 3).
+However, at larger scales, accuracy degrades.
+The decreasing compression ratios with increasing batch size in constant local batch size experiments (Figure 4(a)) suggest this degradation might be due to the loss or underutilization of important smaller singular values during model averaging.
+Similarly, in constant global batch size experiments, increasing compression is coupled with decreasing accuracy (Figures 4(b) and 3).
+This is attributed to larger divergences in smaller singular values as local groups explore the loss landscape independently.
+As this effect strengthens, the divergences between independent groups occur primarily in the smaller, yet important, singular values.
+Averaging them together causes the divergent singular values to shrink relative to the other values.
+Over the course of training, these values can become so small that they are removed, hindering AB training at extreme scales.
+[…]
+Constant Local Batch Size
+ViT B/16 ResNet-50
+Batch Size AB AB - No Groups Traditional DDP AB AB - No Groups Traditional DDP
+2k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.81 $\pm$ 0.08 5.14 $\pm$ 0.19 8.87 $\pm$ 0.00 1.11 $\pm$ 0.00 2.87 $\pm$ 0.01 3.83 $\pm$ 0.01
+Compression 2.52 : 1 2.83 : 1 1 : 1 1.29 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 435.80 $\pm$ 7.36 442.94 $\pm$ 8.52 436.38 $\pm$ 0.35 368.15 $\pm$ 0.80 364.91 $\pm$ 0.53 367.36 $\pm$ 0.36
+Validation Top 1, % 70.24 $\pm$ 0.36 68.31 $\pm$ 0.07 70.73 $\pm$ 0.44 75.67 $\pm$ 0.10 74.42 $\pm$ 0.04 74.43 $\pm$ 0.06
+4k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.83 $\pm$ 0.06 5.72 $\pm$ 0.08 8.87 $\pm$ 0.10 1.04 $\pm$ 0.05 2.67 $\pm$ 0.04 3.85 $\pm$ 0.01
+Compression 2.09 : 1 2.22 : 1 1 : 1 1.39 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 220.85 $\pm$ 0.88 224.35 $\pm$ 1.22 249.03 $\pm$ 31.01 200.69 $\pm$ 10.30 195.46 $\pm$ 1.30 195.19 $\pm$ 0.59
+Validation Top 1, % 70.90 $\pm$ 0.44 67.81 $\pm$ 0.49 69.15 $\pm$ 0.00 74.61 $\pm$ 0.06 73.68 $\pm$ 0.12 73.67 $\pm$ 0.05
+8k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.65 $\pm$ 0.04 5.93 $\pm$ 0.13 8.51 $\pm$ 0.00 0.91 $\pm$ 0.03 2.71 $\pm$ 0.02 3.71 $\pm$ 0.09
+Compression 2.02 : 1 1.98 : 1 1 : 1 1.31 : 1 1.16 : 1 1 : 1
+Time to Train, $\min$ 115.65 $\pm$ 1.37 121.95 $\pm$ 1.95 121.78 $\pm$ 1.02 108.19 $\pm$ 4.94 104.70 $\pm$ 0.52 105.77 $\pm$ 1.69
+Validation Top 1, % 68.92 $\pm$ 0.10 65.70 $\pm$ 0.14 67.20 $\pm$ 0.07 73.66 $\pm$ 0.03 73.04 $\pm$ 0.15 73.02 $\pm$ 0.10
+16k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.54 $\pm$ 0.03 4.91 $\pm$ 0.67 8.25 $\pm$ 0.18 0.87 $\pm$ 0.01 2.56 $\pm$ 0.04 3.63 $\pm$ 0.23
+Compression 1.86 : 1 1.86 : 1 1 : 1 1.19 : 1 1.10 : 1 1 : 1
+Time to Train, $\min$ 64.83 $\pm$ 1.96 79.15 $\pm$ 9.31 63.53 $\pm$ 1.10 58.61 $\pm$ 0.07 59.54 $\pm$ 0.68 59.27 $\pm$ 0.10
+Validation Top 1, % 64.20 $\pm$ 0.01 61.17 $\pm$ 0.33 63.68 $\pm$ 0.09 72.66 $\pm$ 0.04 72.60 $\pm$ 0.10 72.32 $\pm$ 0.03
+32k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.27 $\pm$ 0.41 4.63 $\pm$ 0.51 7.52 $\pm$ 0.36 0.78 $\pm$ 0.01 2.17 $\pm$ 0.04 3.00 $\pm$ 0.06
+Compression 1.80 : 1 1.84 : 1 1 : 1 1.20 : 1 1.12 : 1 1 : 1
+Time to Train, $\min$ 39.53 $\pm$ 5.29 43.36 $\pm$ 3.90 36.56 $\pm$ 0.19 36.91 $\pm$ 0.07 37.48 $\pm$ 0.15 37.91 $\pm$ 0.73
+Validation Top 1, % 55.54 $\pm$ 0.34 53.89 $\pm$ 0.15 58.00 $\pm$ 0.17 70.49 $\pm$ 0.05 71.38 $\pm$ 0.07 71.47 $\pm$ 0.08
+Constant Global Batch Size
+ViT B/16 ResNet-50
+Nodes AB AB - No Groups Traditional DDP AB AB - No Groups Traditional DDP
+4 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.49 $\pm$ 0.49 5.61 $\pm$ 0.16 8.93 $\pm$ 0.09 1.07 $\pm$ 0.01 2.69 $\pm$ 0.03 3.83 $\pm$ 0.01
+Compression 2.09 : 1 2.21 : 1 1 : 1 1.38 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 259.44 $\pm$ 49.68 227.99 $\pm$ 5.29 224.97 $\pm$ 1.79 194.64 $\pm$ 0.91 195.76 $\pm$ 3.49 195.03 $\pm$ 0.74
+Validation Top 1, % 70.68 $\pm$ 0.56 68.66 $\pm$ 0.22 69.26 $\pm$ 0.18 74.75 $\pm$ 0.09 73.65 $\pm$ 0.13 73.74 $\pm$ 0.13
+8 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 4.89 $\pm$ 0.44 10.31 $\pm$ 0.32 13.09 $\pm$ 1.92 1.46 $\pm$ 0.18 3.98 $\pm$ 0.41 4.96 $\pm$ 0.09
+Compression 2.27 : 1 2.19 : 1 1 : 1 1.54 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 123.35 $\pm$ 5.81 124.85 $\pm$ 1.92 153.03 $\pm$ 17.77 135.98 $\pm$ 12.43 131.76 $\pm$ 12.52 139.82 $\pm$ 2.00
+Validation Top 1, % 69.66 $\pm$ 0.44 68.37 $\pm$ 0.15 69.38 $\pm$ 0.14 73.76 $\pm$ 0.19 73.60 $\pm$ 0.02 73.64 $\pm$ 0.09
+16 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 8.17 $\pm$ 0.17 17.14 $\pm$ 0.28 25.00 $\pm$ 0.31 1.91 $\pm$ 0.27 5.57 $\pm$ 0.09 6.09 $\pm$ 0.46
+Compression 2.54 : 1 2.20 : 1 1 : 1 1.72 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 77.68 $\pm$ 9.18 74.26 $\pm$ 0.76 76.95 $\pm$ 0.87 111.46 $\pm$ 12.41 92.52 $\pm$ 1.78 113.00 $\pm$ 7.23
+Validation Top 1, % 69.19 $\pm$ 0.68 68.67 $\pm$ 0.34 68.79 $\pm$ 0.40 72.41 $\pm$ 0.04 73.54 $\pm$ 0.04 73.72 $\pm$ 0.03
+32 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 9.84 $\pm$ 0.07 21.81 $\pm$ 2.66 25.00 $\pm$ 2.55 1.84 $\pm$ 0.42 5.70 $\pm$ 0.03 6.49 $\pm$ 1.85
+Compression 2.91 : 1 2.20 : 1 1 : 1 1.91 : 1 1.27 : 1 1 : 1
+Time to Train, $\min$ 51.10 $\pm$ 0.29 81.41 $\pm$ 19.36 77.99 $\pm$ 18.29 110.09 $\pm$ 28.61 92.37 $\pm$ 3.35 121.55 $\pm$ 31.04
+Validation Top 1, % 66.13 $\pm$ 0.22 68.58 $\pm$ 0.57 68.99 $\pm$ 0.10 70.59 $\pm$ 0.08 73.73 $\pm$ 0.10 73.71 $\pm$ 0.04
+Understanding the interplay between network bandwidth limitations and our method’s potential benefits in real-world environments is crucial.
+The observed accuracy degradation at scale suggests the need to research improved update mechanisms, potentially exploring non-average update rules, mixing matrices, or loss-based weighted averaging schemes.
+VI Conclusion
+Our experimental results demonstrate the significant potential of our AB training method to reduce interconnect traffic by utilizing low-rank representations and independently trained worker subgroups.
+The consistent 70% reduction in network traffic without an increase in training time nor a decrease in the model’s accuracy has the potential to unlock new possibilities within distributed-memory computing environments.
 This could be particularly impactful for scientific research on machines without high-speed interconnects and for exploring even larger, more complex models.
@@
 By addressing the limitations uncovered in our study, we believe the research community can pave the way for even more efficient and scalable training of large-scale neural networks, ultimately furthering scientific discovery across a variety of domains.
-Acknowledgements.
+Acknowledgment
 This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research.